# MultiFinBen: A Multilingual, Multimodal, and Difficulty-Aware Benchmark for Financial LLM Evaluation

### v2

@@ 2.1 Overview
In this section, we present MultiFinBen, the first unified benchmark for evaluating LLMs in the financial domain across diverse modalities, linguistic settings, and tasks (Table  2).
Structured difficulty-aware benchmarking
Rather than aggregating all available datasets from different modalities, linguistics, and tasks, MultiFinBen is organized hierarchically across modality, language, and task category, and for each configuration, we retain a single dataset per difficulty level. This selection policy keeps the benchmark compact and interpretable while enabling seamless integration of new tasks, languages, and modalities  suzgun2022challengingbigbenchtaskschainofthought ; glazer2024frontiermathbenchmarkevaluatingadvanced .
To estimate dataset difficulty, we compute the average standardized performance of two reference models: GPT-4o hurst2024gpt and LLaMA3.1-70B-Instruct dubey2024llama . These models are chosen to span the spectrum of current closed- and open-source capabilities, providing a practical and interpretable signal of how challenging a dataset is across the frontier. Based on this average score, we assign each dataset to one of three difficulty levels: easy (average > 60), medium (20–60), and hard (average < 20). The thresholds are calibrated to reflect distinct performance regimes: scores above 60 typically indicate tasks that both models have largely mastered, scores below 20 indicate consistent failure or poor generalization, and the 20–60 band captures the transitional region where models begin to show meaningful variation (Table 5, Figure 2). This stratification enables layered evaluation, from verifying basic functionality to identifying open challenges and tracking incremental gains.
Once difficulty levels are established, we select one dataset per modality–language–task configuration within each tier. To prioritize datasets that are most informative for evaluation, we rank candidates by the absolute performance gap between the two models. Larger gaps indicate tasks that reveal capability boundaries and model-specific weaknesses, which are especially valuable for benchmarking progress. In cases where multiple datasets exhibit similar contrast, we select the one with lower overall performance to preserve headroom for future improvement.
Linguistics
@@
In MultiFinBen, we are the first to evaluate financial LLMs across three linguistic settings: monolingual, bilingual, and multilingual. In the monolingual setting, models operate within a single language; the bilingual setting involves cross-lingual transferring; and the multilingual setting requires processing across multiple languages simultaneously. To support this, we include five typologically and economically diverse languages, such as English, Chinese, Japanese, Spanish, and Greek, balancing linguistic diversity, writing systems, and global financial regions across North America, East Asia, Southern Europe, and Latin America.
Modalities
In addition to text, we are also the first to include both visual and audio modalities in a unified financial benchmark for evaluating LLMs. In the visual modality, we focus on charts, tabular data, and text embedded in images, reflecting formats commonly found in financial reports and regulatory documents. In the audio modality, we include spoken financial content such as earnings calls, which contain spontaneous speech, domain-specific terminology, and prosodic cues that are essential for understanding intent, sentiment, and narrative context qin2019you ; cao2024ecc . By evaluating models on both modalities, we enable a more rigorous and realistic assessment of their ability to perform in high-stakes, information-rich financial environments.
Tasks categories
Inspired by FinBen xie2024finbenholisticfinancialbenchmark , we organize candidate datasets under a unified taxonomy of seven core financial NLP tasks: Information Extraction (IE) focuses on converting unstructured financial text into structured outputs. Textual Analysis (TA) assesses a model’s ability to interpret sentiment, topic, or tone in financial discourse. Question Answering (QA) evaluates comprehension of financial content through question answering. Text Generation (TG) focuses on producing coherent, informative, and factually accurate financial text. Risk Management (RM) targets detection or analysis of risk-related signals. Forecasting (FO) measures a model’s ability to predict market trends or investor behavior. Decision-Making (DM) simulates complex financial decision processes.
Datasets
To support MultiFinBen, we aggregate datasets from two sources: newly introduced datasets and existing benchmarks. For new datasets, we introduce PolyFiQA-Easy and PolyFiQA-Expert to fill the gap in multilingual financial comprehension and reasoning, and two OCR-based datasets constructed from financial PDFs, the dominant format in global financial communication. For existing resources, we incorporate datasets from FinBen xie2024finbenholisticfinancialbenchmark , AuditWen jiajia-etal-2024-auditwen , FinanceIQ xuanyuan2024 , chABSA  [^2] https://github.com/chakki-works/chABSA-dataset, FLARE-ES 10.1145/3637528.3671554 , Plutus-Ben peng2025plutus , OpenFinLLMs xie2024open , and FinAudio cao2025finaudiobenchmarkaudiolarge , and apply our difficulty-aware mechanism to ensure balanced, multilingual, and multimodal evaluation.
2.2 Text Benchmark
2.2.1 Monolingual
For English, we include thirteen datasets from FinBen xie2024finbenholisticfinancialbenchmark covering all task categories. For IE, SC mariko-etal-2020-financial detects causality in filings and FinRED sharma2022finred extracts relations in news and earnings calls, both evaluated by F1 sokolova2009systematic , while FINER-ORD shah2023finer performs named entity recognition using entity F1 derczynski-2016-complementarity . For TA, Headlines sinha2021impact extracts actionable signals (avg F1) sokolova2009systematic and TSA cortis2017semeval conducts sentiment analysis (accuracy makridakis1993accuracy ). For QA, FinQA chen2021finqa and TATQA zhu2021tat assess numerical reasoning in financial reports and tables, whereas XBRL-Math chen2022convfinqa focus on equation inference (accuracy makridakis1993accuracy ). For TG, EDTSUM xie2023pixiu and ECTSUM mukherjee2022ectsum summarize financial news and earnings calls, respectively, using ROUGE-1 lin-2004-rouge . For RM, CCF feng2024empowering targets fraud detection (Matthews Correlation Coefficient, MCC matthews1975comparison ). For FO, BigData22 soun2022accurate addresses stock movement prediction (MCC matthews1975comparison ). For DM, MSFT yu2024finconsynthesizedllmmultiagent evaluate LLM trading strategies via Sharpe Ratio (SR) sharpe1998sharpe . For Chinese, we include four datasets: RRE jiajia-etal-2024-auditwen for semantic relation extraction (IE); AIE and LNE jiajia-etal-2024-auditwen for audit classification of targets and legal references (TA); and FinanceIQ 10.1145/3583780.3615285 for multi-choice financial QA, all evaluated by classification accuracy makridakis1993accuracy . For Japanese, we include only one public dataset, chabsa kubo2018chabsa , from the Japanese financial benchmark hirano2024construction , which performs sentiment classification in securities filings, evaluated by macro F1 hinojosa2024performance . For Spanish, we include four datasets from FLARE-ES 10.1145/3637528.3671554 : MultiFin jorgensen2023multifin and TSA pan2023evaluation for sentiment and topic classification (TA, accuracy makridakis1993accuracy ); EFPA for multiple-choice QA (accuracy makridakis1993accuracy ); and FNS-2023  10386228 for annual report summarization (TG, ROUGE-1 lin-2004-rouge ). For Greek, we include four datasets from Plutus-ben peng2025plutus , spanning IE, TA, QA, and TG. In IE, GRFinNUM peng2025plutus perform NER for numeral categories (entity F1 derczynski-2016-complementarity ). In TA, GRMultiFin jorgensen-etal-2023-multifin classifies financial headlines (accuracy makridakis1993accuracy ). In QA, GRFinQA peng2025plutus handles domain-specific multiple-choice reasoning (accuracy makridakis1993accuracy ). In TG, GRFNS-2023 10386228 performs financial report summarization (ROUGE-1 lin-2004-rouge ). More details are in Appendix F.
2.2.2 Bilingual
Beyond monolingual benchmarks, we incorporate DOLFIN nakhlé2025dolfindocumentlevelfinancial , a bilingual financial machine translation (MT) dataset under TG task. To align with our language scope, we include the English–Spanish (En–Es) subset from DOLFIN of 1,932 document-level aligned segments. DOLFIN enables assessment of translation coherence and domain fidelity. However, as a parallel resource, it remains limited to single-direction translation between language pairs and does not involve simultaneous multilingual reasoning. Translation quality is measured using Comet-da-22 rei2022cometkiwiistunbabel2022submission (Unbabel/wmt22-comet-da), a reference-based metric optimized for document-level MT.
2.2.3 Multilingual
To address the scarcity of multilingual datasets aligned with our modality-language-task framework, we introduce the first multilingual datasets explicitly designed for multilingual financial reasoning grounded in native-language sources, spanning five economically significant languages (English, Chinese, Japanese, Spanish, and Greek). Our datasets moves beyond existing datasets by directly sourcing native-language financial disclosures and news. This approach preserves real-world complexity and ensures linguistic authenticity, enabling evaluation of LLMs in real-world multilingual financial scenarios. As financial decision-making becomes increasingly cross-border, our datasets offers a critical resource for building globally competent language models.
Task definition
@@ 2.2.3 Multilingual
We define a multilingual financial task that requires models to comprehend over heterogeneous, real-world financial disclosures. Each instance consists of a context $C=\{R,N_{\text{en}},N_{\text{zh}},N_{\text{ja}},N_{\text{es}},N_{\text{el}}\}$, where $R$ denotes a financial report (10-K or 10-Q) and $N_{\text{lang}}$ represents contemporaneous news articles in English, Chinese, Japanese, Spanish, and Greek. These news articles are published near the financial report’s release date and thematically relate to the corresponding report. Given a carefully designed natural language question $q$ and the associated multilingual context $C$, the model must generate an answer $a$ grounded in integrated multilingual information. This task challenges models to perform multilingual cross-document reasoning, integrating disparate pieces of financial signals expressed across diverse languages, as real-world financial analysts must do under time-sensitive, high-stakes conditions.
Data source
@@
We introduce two novel datasets: PolyFiQA-Easy and PolyFiQA-Expert, targeting distinct levels of reasoning complexity. Financial reports are sourced from the SEC, covering 10-K and 10-Q filings  [^3] https://www.sec.gov/. To maintain focus and reduce noise, we extract three core statements from lengthy financial reports: Comprehensive Income, Consolidated Balance Sheets, and Cash Flows  [^4] https://www.sec.gov/. Native-language news articles are collected based on publication date proximity and topic alignment (Appendix H.1). For low-resource languages with sparse coverage, expert-generated news-style texts are validated by native speakers and financial professionals (Appendix H.2).
Expert-in-the-loop data construction
To ensure benchmark fidelity and domain rigor, we adopt an expert-in-the-loop pipeline. Three financial professionals (Appendix H.5) with expertise in economics, business, and accounting oversaw all phases of construction, including news selection, question authoring, guideline development, annotation, and quality control. News articles were meticulously screened for strong alignment with financial reports. Questions were crafted to anchor in real analytical tasks and span two difficulty tiers: easier questions in PolyFiQA-Easy and more complex ones in PolyFiQA-Expert (Appendix H.3). This structure supports fine-grained model assessment across reasoning levels. Rigorous tier-specific annotation guidelines (Appendix H.4) were refined through iterative pilot annotation rounds including tier-specific structures and formatting protocols to promote inter-annotator consistency. In total, 57 expert hours were logged in Label Studio (Appendix H.6), establishing a high quality, reproducible, auditable, and streamlined workflow aligned with best practices in benchmark creation. The raw datasets were further converted into structured instruction datasets with task-specific prompts thoughtfully crafted by financial professionals (Appendix H.9).
Quality validation
Evaluating the quality of free-text generation datasets remains a persistent challenge in both QA and abstractive summarization. To ensure annotation reliability, we adopt a structured scoring framework inspired by prior summarization benchmarks takeshita-etal-2024-aclsum ; gliwa-etal-2019-samsum , evaluating each instance along three key dimensions: Relevance (scored 1–4) captures whether the response includes the key information required to answer the question. Consistency (scored 1–3) measures factual accuracy, especially numerical values. Each question is initially annotated by one expert and independently scored by two additional reviewers using detailed, pilot-refined guidelines (Appendix H.7). Only responses with cumulative scores above 8 are retained in the final dataset. To validate scoring reliability, we report inter-annotator agreement as normalized difference percentage across dimensions. PolyFiQA-Easy and PolyFiQA-Expert achieved average inter-annotator agreements of 89.38%, 91.21%, demonstrating the benchmark’s high quality and scoring consistency.
Evaluation metric
@@
We adopt ROUGE-1 lin-2004-rouge to measure unigram overlap between model predictions and references, offering a proxy for content coverage and factual alignment in multilingual QA tasks.
2.3 Vision Benchmark
For the vision modality in the financial domain, we introduce two novel datasets, EnglishOCR and SpanishOCR, under the IE task, and incorporate an existing dataset, TableBench xie2024open , under the QA task from Open-FinLLMs xie2024open . TableBench xie2024open evaluates multimodal reasoning over tabular images with 450 questions covering comparison and data retrieval, reflecting realistic financial decision-making scenarios.

### v3

@@ 2.1 Overview
Built upon this foundation, 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: systematically organizes datasets along three structural dimensions (i.e., modality, language, and task) to enable comprehensive and interpretable evaluation. Rather than aggregating heterogeneous resources, it introduces a difficulty-aware benchmarking framework that selects representative datasets based on their discriminative power across models, ensuring balanced coverage and meaningful scalability. To address gaps in existing resources, we further design new datasets and tasks that capture underrepresented yet crucial financial scenarios. Together, these principles establish 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: as a unified, rigorous, and extensible framework for assessing financial reasoning across modalities, languages, and task types.
Modalities.
Beyond text, it uniquely integrates visual (charts, tables, images) and audio (earnings calls) modalities into a unified financial benchmark, reflecting common formats of financial dissemination and enabling realistic assessment of multimodal understanding in financial contexts.
Linguistics.
@@
0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: is the first to evaluate LLMs under three linguistic settings (monolingual, bilingual, and multilingual) covering five typologically and economically diverse languages (English, Chinese, Japanese, Spanish, and Greek). This design spans major global financial regions, providing a realistic setting for evaluating cross-lingual financial understanding.
Task Categories.
Inspired by FinBen Xie et al. (2024a), our datasets are organized into seven core financial NLP tasks (Appendix B): Information Extraction (IE), Textual Analysis (TA), Question Answering (QA), Text Generation (TG), Risk Management (RM), Forecasting (FO), and Decision-Making (DM). This organization ensures broad task coverage and supports fine-grained evaluation of reasoning capabilities across diverse financial applications.
2.2 Structured Difficulty-Aware Benchmarking
Existing financial and cross-domain benchmarks often emphasize breadth, aggregating large numbers of datasets across tasks and modalities. While such aggregation maximizes coverage, it obscures the underlying sources of difficulty and makes model progress difficult to interpret. In contrast, 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: adopts a structured, difficulty-aware design that disentangles evaluation along three orthogonal dimensions, i.e., modality, language, and task, enabling controlled comparison and clear attribution of performance to specific reasoning or linguistic factors Suzgun et al. (2022b); Glazer et al. (2024).
To quantify dataset difficulty in a reproducible and model-agnostic manner, we compute the mean standardized performance of two representative large language models, GPT-4o Hurst et al. (2024) and LLaMA-3.1-70B-Instruct Dubey et al. (2024), defined as
$\bar{s}_{d}=\tfrac{1}{2}\big(s_{\text{GPT-4o},d}+s_{\text{LLaMA3.1},d}\big).$ (1)
These models are chosen for their complementary characteristics: GPT-4o provides frontier-level reasoning and instruction alignment, while LLaMA-3.1-70B-Instruct serves as an open, transparent baseline. Their differences in architecture, scale, and training corpus yield a balanced estimate of intrinsic dataset complexity, independent of individual model bias.
We categorize datasets into three tiers, easy ($\bar{s}_{d}>60$), medium ($20\leq\bar{s}_{d}\leq 60$), and hard ($\bar{s}_{d}<20$), corresponding to reliable competence, transitional reasoning, and consistent failure, respectively (Table 5; Figure 7; Appendix D). Within each modality–language–task configuration, we retain the dataset exhibiting the largest inter-model divergence, as such cases most effectively reveal capability boundaries. In the event of ties, the lower-performing dataset is selected to preserve headroom for future progress. This principled selection process yields a compact yet diagnostic benchmark that balances interpretability with challenge, supporting layered evaluation across easy, medium, and hard regimes while maintaining scalability and discriminative power.
2.3 Benchmark Composition
0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: is organized around two complementary sources of datasets, aligning with our three evaluation dimensions: modality, language, and task. (i) Newly introduced resources expand the benchmark’s task coverage, including two multilingual financial QA datasets (PolyFiQA-Easy and PolyFiQA-Expert; Section 2.4.1) and four OCR-based datasets for document-image understanding (EnglishOCR, JapaneseOCR, SpanishOCR, and GreekOCR; Section 2.4.2). (ii) Existing financial benchmarks are integrated through our structured, difficulty-aware selection framework, ensuring balanced representation across text, vision, and audio modalities and across five languages. Together, these resources define the complete task suite of 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, supporting evaluation of financial reasoning, comprehension, and generation across multilingual and multimodal settings.
Text.
In the monolingual setting, 26 datasets across five languages are selected from 68 available candidates following our difficulty calibration (Appendix E). In the bilingual setting, we include the English–Spanish subset of DOLFIN Nakhlé et al. (2025), a document-level financial translation benchmark. As no public datasets exist for multilingual financial reasoning, we introduce two new multilingual QA datasets, PolyFiQA-Easy and PolyFiQA-Expert, designed to evaluate cross-lingual comprehension and reasoning.
Vision.
For the visual modality, we integrate two English QA datasets from Open-FinLLMs Xie et al. (2024c). Among these, TableBench is retained due to its greater inter-model performance variance, offering richer signal for difficulty assessment. We further contribute four newly constructed OCR-based datasets (Section 2.4.2) to evaluate document-level text extraction and structural understanding across languages.
Audio.
For the audio modality, we initially consider five English text-generation datasets from FinAudio Cao et al. (2025b). After applying our structured filtering, two datasets are retained: MDRM Qin and Yang (2019), a medium-tier ASR task featuring short earnings call clips, and FinAudioSum Cao et al. (2025b), a hard-tier summarization task comprising long-form financial recordings (Appendix F).
Together, these resources enable systematic evaluation of financial understanding across text, vision, and audio modalities under consistent difficulty calibration.
2.4 Novel Tasks and Datasets
To close critical gaps in multilingual and visual financial reasoning, we introduce two new task families within 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:: (1) Multilingual Financial Question Answering (i.e., PolyFiQA), the first task for multilingual cross-document reasoning grounded in native financial sources; and (2) Financial OCR, the first OCR task for structured extraction from scanned financial documents. Together, these tasks extend financial evaluation beyond text-based English corpora, bringing cross-lingual reasoning and document-level visual understanding into one unified framework.
2.4.1 Multilingual Financial QA (PolyFiQA)
Unlike prior work relying on translation-based alignment, PolyFiQA draws directly from native-language financial disclosures and contemporaneous news, preserving authentic phrasing, cultural framing, and domain complexity. This design enables realistic evaluation of LLMs in real-world multilingual financial contexts, where decision-making increasingly spans languages and jurisdictions.
Task Definition.
Provided a multilingual context
$C=\{R,N_{\text{en}},N_{\text{zh}},N_{\text{ja}},N_{\text{es}},N_{\text{el}}\},$
@@ 2.4.1 Multilingual Financial QA (PolyFiQA)
where $R$ denotes a financial report (10-K or 10-Q) and $N_{\text{lang}}$ represents contemporaneous news articles in five languages. Given a carefully designed natural language question $q$ and the associated multilingual context $C$, the model must generate an answer $a$ grounded in integrated multilingual information. This task challenges models to perform multilingual cross-document reasoning, mirroring how analysts synthesize signals from diverse linguistic sources under real financial conditions.
Data Construction.
@@
We release two complementary datasets: PolyFiQA-Easy and PolyFiQA-Expert, targeting different reasoning depths. Financial filings are collected from SEC EDGAR  [^2] https://www.sec.gov/search-filings and paired with temporally aligned multilingual news (Appendix H.2). To maintain focus and reduce noise, we extract three core statements from lengthy financial reports: Comprehensive Income, Consolidated Balance Sheets, and Cash Flows. Low-resource cases are supplemented with expert-authored, native-language news verified by bilingual financial professionals (Appendix H.3).
To ensure domain fidelity, we adopt an expert-in-the-loop pipeline (Figure 8): three financial experts curated, annotated, and validated all data, spending over 130 hours in Label Studio (Appendix H.8). Tier-specific guidelines (Appendix H.6) and iterative pilots ensured consistency. Each question was authored and scored for two dimensions, i.e., Relevance (1–4) and Consistency (1–3), by multiple annotators following a structured validation protocol. Only instances with cumulative scores above 5 were retained, yielding inter-annotator agreement above 89% for both tiers. This rigorous process results in a high-quality, auditable dataset for fine-grained evaluation of multilingual financial reasoning. More details can be found in Appendix H.1.
Evaluation Metric.
@@
We adopt ROUGE-1 Lin (2004) to measure unigram overlap between model predictions and references, offering a proxy for content coverage and factual alignment in multilingual QA and TG tasks.